# The Polar Express: Optimal Matrix Sign Methods and Their Application to the Muon Algorithm

### v1

@@ Appendix E Initialization for Matrices with Large Spectral Gaps
In Section 4, we constructed a sequence of polynomials that is adapted to the range of the singular values $[\ell,u]$. Assuming nothing else about the input, these polynomials are optimal because they provide a good approximation to $1$ across the entire interval. However, in many applications, the spectrum has large gaps; that is, there are several large outlying singular values that are well-separated from the rest, . For these matrices, it is not necessary for the polynomial to be accurate on the entire interval $[\ell,u]$, only on the range of the small singular values plus a few large isolated points. In this section, we take advantage of this structure to accelerate our method by preprocessing the matrix to eliminate the largest singular values.
@@
$\sigma_{i}^{2}=\|\bm{M}\|_{\text{F}}^{2}-\sum\limits_{j=1,j\neq i}^{n}\sigma_{j}^{2}\leq\|\bm{M}\|_{\text{F}}^{2}-\sum\limits_{j=1,j\neq i}^{k}\sigma_{j}^{2}\leq\|\bm{M}\|_{\text{F}}^{2}-\sum\limits_{j=1,j\neq i}^{k}\tilde{\sigma}_{j}^{2}$ (21)
@@
$\sigma_{i}\in\left[\tilde{\sigma}_{i},\,\,\sqrt{\|\bm{M}\|_{\text{F}}^{2}-\sum\limits_{j=1,j\neq i}^{k}\tilde{\sigma}_{j}^{2}}\right]$
@@
Because $p(0)=0$ and $p$ has at most one local extremum on $\mathbb{R}_{\geq 0}$, these conditions immediately guarantee that $p$ is concave-increasing on $[0,\sqrt{1-z^{2}}]$, so it must lie above the line $x\mapsto x/\sqrt{1-z^{2}}$. Furthermore, $p$ is decreasing on $[\sigma_{1},1]$, so it maps $\sigma_{1}\in[z,1]$ to $[p(1),1]$. By minimizing $p(1)$ over all valid $z$s (that is, over the interval $[1/\sqrt{2},1]$), one can further show that $p(1)>1/\sqrt{2}$, so $\sigma_{1}$ cannot be decreased very much by applying $p$. Thus, the largest singular value of $p(\bm{M})$ is still at most $1$, while the smaller singular values have increased by a potentially large factor of $1/\sqrt{1-z^{2}}$. When there is a large outlying singular value, $z$ is close to $1$ and this initialization scheme makes much more progress than a standard iteration of PolarExpress would have.
In Figure 7, we demonstrate the benefit of using the $p$ given by (22) on a synthetic matrix whose spectrum follows a power law decay. That is, $\sigma_{j}(\bm{M})=j^{-5}$. For both Newton-Schulz and Polar Express performing the extra spectrum-aware initialization step described in this section leads to significant speedups in convergence, even if we use one fewer iteration of the baseline method
@@ Appendix F Fast Polynomial Iteration for Rectangular Matrices
In this section, we describe a simple method for applying an iterative polynomial method to a rectangular matrix. For matrices with a large aspect ratio, this method yields significant computational savings. We emphasize that this method is applicable to any computation of the form $(p_{T}\circ\cdots\circ p_{1})(\bm{X})$, where each $p_{t}$ is an odd polynomial.
As a preliminary, we first describe the baseline approach. Let $\bm{X}\in\mathbb{R}^{m\times n}$ with $m=\alpha n\geq n$, where $\alpha\geq 1$ is called the aspect ratio. Any odd polynomial $p$ of degree $d=2q+1$ can be represented as $p(x)=xh(x^{2})$, where $h$ is a polynomial of degree $q$. Thus, $p(\bm{X})=\bm{X}h(\bm{X}^{\top}\bm{X})$. Furthermore, $h$ can be written in a factored form called Horner’s rule to reduce the number of multiplications. For instance, if $h(y)=a+by+cy^{2}+dy^{3}$, Horner’s rule gives $h(y)=a+y\left(b+y\left(c+dy\right)\right)$. For a matrix, $h(\bm{Y})=a\bm{I}+\bm{Y}\left(b\bm{I}+\bm{Y}\left(c\bm{I}+d\bm{Y}\right)\right)$. Thus for $\bm{Y}\in\mathbb{R}^{n\times n}$, computing $h(\bm{Y})$ costs about $\left(\deg(h)-1\right)\cdot n^{3}$ operations, and computing $p(\bm{X})=\bm{X}h(\bm{X}^{\top}\bm{X})$ costs $2mn^{2}+\left(\frac{d-1}{2}-1\right)\cdot n^{3}=\left(\frac{d-3}{2}+2\alpha\right)\cdot n^{3}$ operations. This process could be repeated for each iteration $p_{1},\ldots,p_{T}$. Notice that if we instead computed $h(\bm{X}\bm{X}^{\top})\bm{X}$, the result would be the same but the cost would be higher.
@@
Figure 8 shows that using Algorithm 3 can dramatically improve runtime on the GPU when the aspect ratio is large enough. As expected, using Algorithm 3 for many iterations significantly reduces the dependence of the runtime on the aspect ratio. Running six iterations of a degree-5 polynomial method when $\alpha=4$ (as with the linear transformations in each MLP block of a transformer) we obtain almost a 2x speedup, and when $\alpha=32$, we obtain nearly a 10x speedup. If we restart every three iterations, the trend is the same but the runtimes savings are less.
Figure 8: Effects of using Algorithm 3 on runtime on a GPU. We run $T=6$ iterations of a degree-5 polynomial method on matrices with various dimensions $n$ and aspect ratios $\alpha$. Restart interval 6 is Algorithm 3, restart interval 1 is equivalent to not using Algorithm 3, and restart interval 3 is an intermediate method that calls Algorithm 3 once to do the first three iterations and again to do the last three iterations for greater stability. When $\alpha\gg 1$, increasing the restart interval makes the runtime many times faster.
@@ Appendix G Code for Constructing Polynomials of Polar Express
The following code gives a Python implementation of the offline stage of Algorithm 1. This code was used to construct the coefficients of the polynomials given in (17), which in turn were used in our Muon experiments (Section 5.2). It uses $\ell=10^{-3}$ and $u=1$ by default. It incorporates Algorithm 2 and the numerical tweaks described in Section 4.4
@@
# Safety factor:
a /= 1.01; b /= 1.01**3; c /= 1.01**5

### v2

@@ Appendix E Initialization for Matrices with Large Spectral Gaps
In Section 4, we constructed a sequence of polynomials that is adapted to the range of the singular values $[\ell,u]$. Assuming nothing else about the input, these polynomials are optimal because they provide a good approximation to $1$ across the entire interval. However, in many applications, the spectrum has large gaps; that is, there are several large outlying singular values that are well-separated from the rest. For these matrices, it is not necessary for the polynomial to be accurate on the entire interval $[\ell,u]$, only on the range of the small singular values plus a few other isolated points. In this section, we take advantage of this structure to accelerate our method by preprocessing the matrix to eliminate the largest singular values.
@@
$\sigma_{i}^{2}=\|\bm{M}\|_{\text{F}}^{2}-\sum\limits_{\begin{subarray}{c}j=1\\ j\neq i\end{subarray}}^{n}\sigma_{j}^{2}\leq\|\bm{M}\|_{\text{F}}^{2}-\sum\limits_{\begin{subarray}{c}j=1\\ j\neq i\end{subarray}}^{k}\sigma_{j}^{2}\leq\|\bm{M}\|_{\text{F}}^{2}-\sum\limits_{\begin{subarray}{c}j=1\\ j\neq i\end{subarray}}^{k}\tilde{\sigma}_{j}^{2}$ (19)
@@
$\sigma_{i}\in\left[\tilde{\sigma}_{i},\,\,\sqrt{\|\bm{M}\|_{\text{F}}^{2}-\sum\limits_{\begin{subarray}{c}j=1\\ j\neq i\end{subarray}}^{k}\tilde{\sigma}_{j}^{2}}\right]$
@@
Because $p(0)=0$ and $p$ has at most one local extremum on $\mathbb{R}_{\geq 0}$, these conditions immediately guarantee that $p$ is concave-increasing on $[0,\sqrt{1-z^{2}}]$, so it must lie above the line $x\mapsto x/\sqrt{1-z^{2}}$. Furthermore, $p$ is decreasing on $[\sigma_{1},1]$, so it maps $\sigma_{1}\in[z,1]$ to $[p(1),1]$. By minimizing $p(1)$ over all valid $z$ (that is, over the interval $z\in[1/\sqrt{2},1]$), one can further show that $p(1)>1/\sqrt{2}$, so $\sigma_{1}$ cannot be decreased very much by applying $p$. Thus, the largest singular value of $p(\bm{M})$ is still at most $1$, while the smaller singular values have increased by a potentially large factor of $1/\sqrt{1-z^{2}}$. When there is a large outlying singular value, $z$ is close to $1$ and this initialization scheme makes much more progress than a standard iteration of PolarExpress would have.
In Figure 7, we demonstrate the benefit of using the $p$ given by (20) on a synthetic matrix whose spectrum follows a power law decay. That is, $\sigma_{j}(\bm{M})=j^{-5}$, so this matrix has a large outlying singular value $\sigma_{1}\gg\sigma_{2}$. Applying (20) costs almost as much as performing an iteration of a degree-5 polynomial method, so for fair comparison, we count it as an additional iteration in this plot. For both Newton-Schulz and Polar Express, performing the extra spectrum-aware initialization step described in this section leads to significant speedups in convergence.
@@ Appendix F Fast Polynomial Iteration for Rectangular Matrices
In this section, we describe a simple method for applying an iterative polynomial method to a rectangular matrix. For matrices with a large aspect ratio, this method yields significant computational savings. We emphasize that this method is applicable to any computation of the form $(p_{T}\circ\cdots\circ p_{1})(\bm{X})$, where each $p_{t}$ is an odd polynomial. Thus, it can be used to apply Newton-Schulz or Jordan’s polynomials in addition to our own.
As a preliminary, we first describe the baseline approach. Let $\bm{X}\in\mathbb{R}^{m\times n}$ with $m\geq n$, where $\alpha:=m/n\geq 1$ is called the aspect ratio. Any odd polynomial $p$ of degree $d=2q+1$ can be represented as $p(x)=xh(x^{2})$, where $h$ is a polynomial of degree $q$. Thus, $p(\bm{X})=\bm{X}h(\bm{X}^{\top}\bm{X})$. Furthermore, $h$ can be written in a factored form called Horner’s rule to reduce the number of multiplications. For instance, if $h(y)=a+by+cy^{2}+dy^{3}$, Horner’s rule gives $h(y)=a+y\left(b+y\left(c+dy\right)\right)$. For a matrix, $h(\bm{Y})=a\bm{I}+\bm{Y}\left(b\bm{I}+\bm{Y}\left(c\bm{I}+d\bm{Y}\right)\right)$. Thus for $\bm{Y}\in\mathbb{R}^{n\times n}$, computing $h(\bm{Y})$ costs about $\left(\deg(h)-1\right)\cdot n^{3}$ operations, and computing $p(\bm{X})=\bm{X}h(\bm{X}^{\top}\bm{X})$ costs $2mn^{2}+\left(\frac{d-1}{2}-1\right)\cdot n^{3}=\left(\frac{d-3}{2}+2\alpha\right)\cdot n^{3}$ operations. This process could be repeated for each iteration $p_{1},\ldots,p_{T}$. Notice that if we instead computed $h(\bm{X}\bm{X}^{\top})\bm{X}$, the result would be the same but the cost would be higher.
@@
Figure 8 shows that using Algorithm 4 can dramatically improve runtime on the GPU when the aspect ratio is large enough. As expected, using Algorithm 4 for many iterations significantly reduces the dependence of the runtime on the aspect ratio. Running six iterations of a degree-5 polynomial method when $\alpha=4$ (as with the linear transformations in each MLP block of a transformer) we obtain almost a 2x speedup, and when $\alpha=32$, we obtain nearly a 10x speedup. If we restart every three iterations, the trend is the same but the runtime savings are somewhat smaller.
Figure 8: Effects of using Algorithm 4 on runtime on a GPU. We run $T=6$ iterations of a degree-5 polynomial method on matrices with various dimensions $n$ and aspect ratios $\alpha$. Restart interval $=6$ is Algorithm 4, restart interval $=1$ is equivalent to the baseline (that is, not using Algorithm 4), and restart interval $=3$ is an intermediate method that calls Algorithm 4 once to do the first three iterations and again to do the last three iterations for greater stability. When $\alpha\gg 1$, increasing the restart interval significantly reduces the runtime.
@@ Appendix G Code for Constructing Polynomials of Polar Express
The following code gives a Python implementation of the offline stage of Algorithm 2. This code was used to construct the coefficients of the polynomials given in (LABEL:eq:coeffs), which in turn were used in our Muon experiments (Section 5.2). It uses $\ell=10^{-3}$ and $u=1$ by default. It incorporates Algorithm 3 and the numerical tweaks described in Section 4.4
@@
# Optionally incorporate safety factor here:
# a /= 1.01; b /= 1.01**3; c /= 1.01**5